# Comparing diversity, negativity, and stereotypes in Chinese-language AI technologies: a case study on Baidu, Ernie and Qwen

### v1

@@ 4.3 Measuring stereotypes in Ernie and Qwen
Following Choenni et al., (2021), we attempt to measure the prevalence of stereotypes in Ernie and Qwen by analyzing the overlap in completions between the two LLMs and Baidu, based on the assumption that candidate words suggested by the search engine result from stereotypical views expressed by individuals in their online queries. We remark that we cannot reproduce the original approach, which computed the tipicality of views encoded in LLMs about different groups Kurita et al., (2019) based on their internal tokens’ probability as we do not have access to them in the chosen commercial LLMs under analysis.
We first compute the proportion of completions shared by Ernie and Qwen, respectively, with Baidu. Overall, Ernie and Qwen share 26.52% and 27.81% of their completions with Baidu (Figure 12, right panel), meaning that 1 out of 3 candidate words generated by the LLMs to describe different social groups coincide with the views embedded in Chinese online search queries. We regard these overlapping completions as stereotypical. In the left panel of Figure 12 we report the breakdown of the stereotypical attributes across categories, reporting the proportion of attributes associated with a negative sentiment. We can observe that the two LLMs exhibit similar proportions of overlapping output, with Age, Gender, Nationality, Race, and Region being the categories with the largest prevalence of stereotypes ($>20\%$), probably because they have a much larger number of results compared to other categories (cf. Figure 2). Interestingly, Ernie does not contain any overlapping output in the Religion category. We then compute the relative proportion of negative and derogatory stereotypes from Baidu present overall in the output generated by the two LLMs, finding that Ernie has a smaller proportion of negative stereotypes (8.96%) while for Qwen this value is higher (12.43%).
[…]
Figure 12: Proportion of stereotypical views in LLMs’ completions based on Baidu’s completions, weighted by occurrences, across categories (left) and overall (right). The proportion of negative words is written and highlighted with hatched bars.
@@ 4.4 Measuring agreement of LLMs with generated output
We compute the proportion of statements about social groups on which Ernie and Qwen agree based on the sets of unique completions generated by the three models and show them respectively in Figures 1314. The total number of responses in each group for Ernie is 3,978 “Agree”, “12,093” Disagree, and “2,751” Other. For Qwen, there are 5,870 “Agree”, 12,541 “Disagree”, and 411 “Other”.
We can observe that Ernie tends to agree much more on output generated by the model itself (37.6%) and less with the completions of Baidu (19.24%) or Qwen (15.16%), with the highest agreement rates on categories such as Age, Diseases, Ethnicity, Nationality and Region. The proportion of negative views on which the model exhibits agreement is very low ($<3\%$) across the three groups of completions.
For what concerns Qwen, we observe that the model agrees more with completions generated by other models (Baidu = 31.57%, Ernie = 45.09%) than itself (21.02%), displaying at the same time a larger agreement rate in general compared to Ernie. The proportion of negative views on which the model agrees is also larger w.r.t to the other LLM, with values in the range 2-6%.
Figure 13: (left) Proportion of unique completions for which Ernie expresses agreement across categories and sources of completions, out of all statements where the model either agrees or disagrees. (right) Overall proportion of unique completions for which Ernie expresses agreement across sources of completions, out of all statements where the model either agrees or disagrees.
Figure 14: (left) Proportion of unique completions for which Qwen expresses agreement across categories and sources of completions, out of all statements where the model either agrees or disagrees. (right) Overall proportion of unique completions for which Qwen expresses agreement across sources of completions, out of all statements where the model either agrees or disagrees.
@@ 5.1 Contributions
Language models and online search engines can unintentionally perpetuate stereotypes if they consistently generate biased descriptions for certain social groups. Observing a variety of generated descriptions suggests that the language model captures the underlying diversity more accurately, avoiding oversimplified or biased portrayals of social groups, thus better representing the richness of human experiences. Moreover, if a model consistently generates negative or derogatory descriptions, it can perpetuate harmful biases and contribute to societal discrimination, shaping or reinforcing negative perceptions about those social groups. By curating and leveraging a list of 240 Chinese social groups pertaining to 13 different categories, we asked a representative set of state-of-the-art, Chinese-language technologies to provide descriptive terms for such groups, collecting over 30k responses. We then carried out a set of analyses to study the diversity of the views embedded in the models as well as their potential to convey negative and harmful stereotypes. Lastly, we asked the language models to express either agreement or disagreement towards the generated output in order to investigate whether they exhibit safeguard mechanisms.
@@
We show that the LLMs exhibit a much larger diversity of views about social groups compared to Baidu search engine completions, although the candidate words returned to describe groups within the same category yield a much larger overlap compared to groups belonging to different categories. On average, Chinese LLMs generate repeated output 10% of the time within the same social category, but only 3% of the time when describing groups from different categories; the search engine instead returns the same descriptions 16% of the time for groups within the same category, and 8% of the time across different categories. Baidu and Qwen exhibit concerning levels of potentially offensive generated content (1 out of 3 candidate words has a negative sentiment) compared to Ernie, which appears much safer (only approximately 1 out of 10 candidate words is negative). The agreement on negative views about the same social group embedded in the models is moderately high, with significant positive correlations in the range 0.28-0.47, most likely because they were trained on similar datasets that encompasses such social biases. Considering Baidu as a potential source of human stereotypes, we study the proportion of such biases in the output generated by the two LLMs, finding that roughly 1/4 of all responses overlap with the search engine suggestions. Qwen also exhibits a larger prevalence of negative and derogatory stereotypes than Ernie, although considering the overlaps, the difference between the two LLMs is less marked. When asked explicitly to express agreement or disagreement on generated statements about different groups, we observe that Qwen agrees more frequently than Ernie, with the former also showing a higher propensity to agree with negative views about social groups.
@@ 5.2 Implications
Our study reveals that while Large Language Models show a broader diversity of perspectives on social groups compared to traditional search engines like Baidu, they still exhibit a significant amount of repetition within specific social categories. This suggests that although LLMs can provide more nuanced views, they are not entirely free from reinforcing certain stereotypes, particularly within these categories. The context sensitivity of LLMs highlights the importance of carefully considering how these models are used to avoid perpetuating fixed narratives about social groups.
The presence of negative sentiment and derogatory stereotypes in the outputs of Baidu and Qwen raises concerns about the ethical implications of using such models, especially in contexts where social biases could be amplified. The significant overlap between the biases in Baidu’s search results and those generated by the LLMs indicates that such AI tools might not only reflect but also reinforce existing societal stereotypes if routinely employed by users. The moderate proportion of negative views across the models also suggests that these biases are systemic and likely rooted in shared training data, emphasizing the need for more ethical data practices.
@@ 5.3 Limitations and future work
Our study, while covering a broad range of social groups in Chinese society, may not capture all relevant social terms or cultural nuances. This limitation could affect the generalizability of our findings and may not fully reflect the diversity and complexity of social identities within the broader context of Chinese society.
Our use of six templates to gather social group attributes may not fully capture all potential characteristics. This limitation could result in an incomplete representation of the social groups encoded in the LLMs and may overlook important nuances in their description.
To study the negativity of output generated by the three technologies, we relied on Aliyun’s sentiment analysis tool, which may have impacted sentiment labelling. Different tools might yield varying results, and this reliance could affect the precision of our findings regarding negative content.
In another analysis, we used Baidu as a source to study societal stereotypes, assuming its outputs reflect societal biases. However, search engine results are influenced by various factors, which may not always accurately represent stereotypes, and our study does not explore whether these stereotypes are more distinguishable in comparison to those generated by LLMs.
Our work is also limited to two Chinese LLMs, Ernie and Qwen, and does not include other models or those from Western contexts.
@@
Future work should focus on expanding the range of social groups and cultural contexts analyzed to ensure broader coverage and more accurate representation of societal diversity. Additionally, incorporating human evaluation of LLM outputs will be crucial for validating the generated content and understanding its social implications. Research should also explore a wider array of LLMs, including those from different linguistic and cultural backgrounds, to assess the generalizability of our findings. Finally, developing and implementing effective bias mitigation strategies will be essential to address the ethical challenges posed by LLMs, ensuring their responsible use in various applications.

### v3

@@ 4.3 Measuring stereotypes in Ernie and Qwen
Following Choenni et al., (2021), we attempt to measure the prevalence of stereotypes in Ernie and Qwen by analyzing the overlap in completions between the two LLMs and Baidu, based on the assumption that candidate words suggested by the search engine result from stereotypical views expressed by individuals in their online queries. We validated this assumption by manually investigating 100 completions in the overlapping data – the grey region highlighted in Figure 4, left panel – and found that they all referred to stereotypes. Examples include Qwen describing “{CJK*}UTF8gbsn俄罗斯人” (Russians) as “{CJK*}UTF8gbsn爱喝酒” (heavy drinkers) and “{CJK*}UTF8gbsn美国白人” (white Americans) as “{CJK*}UTF8gbsn自以为是” (arrogant), and Ernie labelling “{CJK*}UTF8gbsn比利时人” (Belgians) as “{CJK*}UTF8gbsn懒” (lazy) and “{CJK*}UTF8gbsn为什么日裔这么” (people of Japanese descent) as “{CJK*}UTF8gbsn有礼貌” (polite).
We remark that we cannot reproduce the original approach used in Choenni et al., (2021), which computed the typicality of views encoded in LLMs about different groups based on their internal tokens’ probability Kurita et al., (2019), neither we can employ other probability- or embedding-based approaches reported in the literature Gallegos et al., (2024), as we do not have access to this information in the LLMs under examination.
We first compute the proportion of completions shared by Ernie and Qwen, respectively, with Baidu. Overall, Ernie and Qwen share 26.52% and 27.81% of their completions with Baidu (Figure 12, right panel) meaning that 1 out of 3 candidate words generated by the LLMs to describe different social groups coincide with the views embedded in Chinese online search queries. In the left panel of Figure 12 we report the breakdown of the stereotypical attributes across categories, reporting the proportion of attributes associated with a negative sentiment. We can observe that the two LLMs exhibit similar proportions of overlapping output, with Age, Gender, Nationality, Race, and Region being the categories with the largest prevalence of stereotypes ($>20\%$). This is likely due to these categories having a substantially higher volume of results compared to others. (cf. Figure 2). Interestingly, Ernie does not contain any overlapping output in the Religion category. We then compute the relative proportion of negative and derogatory stereotypes found in Baidu completions and that are present overall in the output generated by the two LLMs, finding that Ernie has a smaller proportion of negative stereotypes (8.96%) compared to Qwen, which exhibits a higher proportion (12.43%).
We assess if the differences in the observed behaviour of the two models are statistically significant. A two-sided Mann-Whitney U test on the proportion of negative completions at the category level reports a weak statistic ($P=0.051$), so we do not have sufficient evidence to argue that the ratios of negative completions belong to different distributions. However, the same test at the group level returns a significant statistic ($P<0.001$). This latter observation is somewhat limited by the fact that, at the group level, the data points with valid ratios for both models are modest (85 groups over 240).
Figure 12: Proportion of stereotypical views in LLMs’ completions based on Baidu’s completions, weighted by occurrences, across categories (left) and overall (right). The proportion of negative words is highlighted with hatched bars and reported as a percentage.
@@ 4.4 Measuring agreement of LLMs with generated output
We compute the proportion of statements about social groups on which Ernie and Qwen agree, based on the sets of unique completions generated by the three models, and show them respectively in Figures 1314. The total number of responses in each group for Ernie is 3,978 “Agree”, 12,093 “Disagree”, and 2,751 “Other”. For Qwen, there are 5,870 “Agree”, 12,541 “Disagree”, and 411 “Other”.
We can observe that Ernie tends to agree much more on the output generated by the model itself (37.6%) and less with the completions of Baidu (19.24%) or Qwen (15.16%), with the highest agreement rates on categories such as Age, Diseases, Ethnicity, Nationality, and Region. The proportion of negative views on which the model exhibits agreement is very low ($<3\%$) across the three groups of completions. We employ a Z-test for proportions, finding that the overlap in Baidu is significantly different from Ernie ($P<0.001$) but not Qwen ($P<0.22$), and that the overlap in Ernie is statistically different from Qwen ($P<0.001$).
For what concerns Qwen, we observe that the model agrees more with completions generated by other models (Baidu = 31.57%, Ernie = 45.09%) than itself (21.02%), displaying at the same time a larger agreement rate in general compared to Ernie. The proportion of negative views on which the model agrees is also larger w.r.t to the other LLM, with values in the range 2-6%. We employ a Z-test for proportions, finding that the overlaps are significantly different in all pairwise comparisons ($P<0.001$).
Figure 13: (left) Proportion of unique completions for which Ernie expresses agreement across categories and sources of completions, out of all statements where the model either agrees or disagrees. (right) Overall proportion of unique completions for which Ernie expresses agreement across sources of completions, out of all statements where the model either agrees or disagrees. A Z-test for proportions finds that the overlap in Baidu is significantly different from Ernie ($P<0.001$) but not from Qwen ($P<0.22$). The proportion of overlap in Ernie is statistically different from Qwen ($P<0.001$).
Figure 14: (left) Proportion of unique completions for which Qwen expresses agreement across categories and sources of completions, out of all statements where the model either agrees or disagrees. (right) Overall proportion of unique completions for which Qwen expresses agreement across sources of completions, out of all statements where the model either agrees or disagrees. A Z-test for proportions finds that the three proportions are statistically different ($P<0.001$).
@@ 5.1 Contributions
Language models and online search engines may inadvertently reinforce stereotypes when they systematically associate biased descriptions with specific social groups. Observing a variety of generated descriptions suggests that the language model captures the underlying diversity more accurately, avoiding oversimplified or biased portrayals of social groups, thus better representing the richness of human experiences Kirk et al., (2023). Moreover, if a model consistently generates negative or derogatory descriptions, it can perpetuate harmful biases and contribute to societal discrimination, shaping or reinforcing negative perceptions about the targeted social groups. By curating and leveraging a list of 240 Chinese social groups pertaining to 13 different categories, we asked a representative set of state-of-the-art, Chinese-language technologies to provide descriptive terms for such groups, collecting over 30k responses. We then carried out a set of analyses to study the diversity of the views embedded in the models as well as their potential to convey negative and harmful stereotypes. Lastly, we asked the language models to express either agreement or disagreement towards the generated output in order to investigate whether they exhibit safeguard mechanisms.
@@
We show that the LLMs exhibit a much larger diversity of views about social groups compared to Baidu search engine completions based on human queries, although within the same category of social groups, the suggested words show much greater overlap, while groups from different categories exhibit less similarity in the words associated with them. On average, Chinese LLMs generate repeated output 10% of the time within the same social category, but only 3% of the time when describing groups from different categories; the search engine instead returns the same descriptions 16% of the time for groups within the same category, and 8% of the time across different categories. Baidu and Qwen exhibit concerning levels of potentially offensive generated content (1 out of 3 candidate words has a negative sentiment) compared to Ernie, which appears much safer in comparison (only approximately 1 out of 10 candidate words is negative). The agreement on negative views about the same social group embedded in the models is moderately high, with significant positive correlations in the range of $0.28-0.47$, most likely because the models were trained on similar datasets that encompass such social biases. Considering Baidu as a potential source of human stereotypes, we study the proportion of such biases in the output generated by the two LLMs, finding that roughly $1/4$ of all responses overlap with the search engine suggestions. Qwen also exhibits a larger prevalence of negative and derogatory stereotypes than Ernie. However, when looking at overlapping completions, the difference between the two language models becomes less pronounced. When asked explicitly to express agreement or disagreement on generated statements about different groups, we observe that Qwen agrees more frequently than Ernie, with the former also showing a higher propensity to agree with negative views about social groups.
@@ 5.2 Implications
Our study reveals that while Large Language Models show a broader diversity of perspectives on social groups compared to traditional search engines like Baidu, they still exhibit a significant amount of repetition within specific social categories. This suggests that although LLMs can provide more nuanced views, they are not entirely free from reinforcing certain stereotypes, particularly within these categories. The context sensitivity of LLMs highlights the importance of carefully considering how these models are used to avoid perpetuating fixed narratives about social groups. Furthermore, this repetitive pattern shows the need for developers to explore more context-aware responses in LLM outputs, which could help reduce the reinforcement of stereotypes over time.
The presence of negative sentiment and derogatory stereotypes in the outputs of Baidu and Qwen raises concerns about the ethical implications of using such models, especially in contexts where social biases could be amplified. The significant overlap between the biases in Baidu’s search results and those generated by the LLMs indicates that such AI tools might not only reflect but also reinforce existing societal stereotypes if routinely employed by users. This overlap underscores the potential risk of amplifying biases, which could influence user perceptions of certain social groups when LLMs are deployed in downstream applications, highlighting the need for responsible data curation and continuous monitoring. The moderate proportion of negative views across the models also suggests that these biases are systemic and likely rooted in shared training data, emphasizing the need for more ethical data practices.
@@ 5.3 Limitations and future work
Our study, while covering a broad range of social groups in Chinese society, may not capture all relevant social terms or cultural nuances. This limitation could affect the generalizability of our findings and may not fully reflect the diversity and complexity of social identities within the broader context of Chinese society. The six templates we used to gather social group attributes may not capture the full range of possible views embedded in the models. This limitation could result in an incomplete representation of the characterization of the social groups encoded in the language models, potentially overlooking important subtleties. Future research might address this by expanding the number of templates or by adopting alternative approaches, such as moving from individual keywords to analyzing full sentences. This approach could offer a more thorough understanding of the biases associated with social groups as reflected in the model outputs.
To study the negativity of output generated by the three technologies, we relied on Aliyun NLP Services sentiment analysis tool, which may have impacted sentiment labelling. Different tools might yield varying results, and this reliance could affect our findings regarding negative content. Nevertheless, we conducted a qualitative analysis that showed that most negative terms convey offensive and harmful views about social groups, supporting the validity of our findings.
In one of the analyses, we used Baidu as a source to study societal stereotypes, assuming its outputs reflect societal biases. However, search engine results are influenced by various factors, which may not always accurately represent stereotypes, and our study does not explore whether these stereotypes are more distinguishable in comparison to those generated by LLMs. However, we carried out a qualitative analysis to investigate the robustness of this approach, which yielded results consistent with this assumption.
Our work is limited to two Chinese LLMs, Ernie and Qwen, and does not include other models, including those from Western contexts. Additionally, since both language models are closed-source, we lack access to crucial internal information, such as model parameters, probability distributions, or training data specifics. This restriction makes it difficult to fully understand the factors contributing to the prevalence or intensity of negative content in their outputs.
@@
Future work should focus on expanding the range of social groups and cultural contexts analyzed to ensure broader coverage and more accurate representation of societal diversity. Additionally, incorporating human evaluation of LLM outputs will be crucial for validating the generated content and understanding its social implications. Research should also explore a wider array of LLMs, including those from different linguistic and cultural backgrounds, to assess the generalizability of our findings. Finally, developing and implementing effective bias mitigation strategies will be essential to address the ethical challenges posed by LLMs, promoting their responsible use across various application domains.
Acknowledgments
During the preparation of this work, the authors used OpenAI ChatGPT in order to proof-check the grammar of some paragraphs and refine the language. After using this tool/service, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication.